OpenStudy (anonymous):

Make a conjecture about the pattern of the given data. Find the sum of the 7th row. 3 3 + 6 + 3 3 + 6 + 9 + 6 + 3 3 + 6 + . . . + ( 3n − 3) + ( 3n ) + ( 3n − 3) + . . . + 3

OpenStudy (anonymous):

So for the first row its \(3(1)=3(1^2)=3\) Second row its\( 3(1+2+1)=3(4)=3(2^2)=12\) Third row is \(3(1+2+3+2+1)=3(9)=3(3^2)=27\) Fourth row is \(3(1+2+3+4+3+2+1)=3(16)=3(4^2)=48\) Fifth Row is \(3(1+2+3+4+5+4+3+2+1)=3(25)=3(5^2)=75\) Do you see a pattern?

OpenStudy (anonymous):

I think I get it now :) So this means that row 6 would be 3(6²) = 108 and 3(7²) = 147.

OpenStudy (anonymous):

YAAAAAA!!!!!

OpenStudy (anonymous):

You made this very easy to understand, thanks a lot :)

OpenStudy (anonymous):

np :)